Find the function (fo h) and simplify. f(x) = 7x + 10, h(x) = Vx+6 (fo h)(x) = 12 (Simplify your answer.)
Answer #1

Given f () = 70+10 bex) = Wato ffon) (0) = f (h (ay) f(wato) ple-rate my answee 7 Sa+6 +10 는 thank you